OpenStudy (anonymous):

Limit x tends to 0. (2^x - 3^x)*x /(1-cos3x)

OpenStudy (welshfella):

for x = 0 we get an indeterminate value 0/0 so we can apply l'hopital' s rule

OpenStudy (anonymous):

L'H'S twice seems like something I would try.

OpenStudy (anonymous):

the second derivative of 1-cos(3x) is 1-cos(3x) 3sin(3x) 9cos(3x) the second derivative of 2^x-3^x is 2^x (ln 2)^2 - 3^x (ln 3)^2 so Lim 2^x (ln 2)^2 - 3^x (ln 3)^2 x -> 0 --------------------------- 9cos(3x) and then you can plug in 0 for x

OpenStudy (anonymous):

no wrong

OpenStudy (anonymous):

because 2^x (ln 2) - 3^x (ln 3) at x=0 is not equal to 0

OpenStudy (anonymous):

then we couldn't use LHS twice, sorry

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{x(2^x-3^x)}{1-\cos(3x)}\]wait, was it this, I made a mistake on interpretation of the problem.

OpenStudy (welshfella):

why did you think it necessary to apply LH twice?

OpenStudy (welshfella):

oh I see why - the denominator would be 3 sin3x which would give a denominator of 0

OpenStudy (anonymous):

Yes fB you right with the problem

OpenStudy (welshfella):

I got the following with applying LH twice -2 * 3^x ln3 - 3^x* x ln^2 (3) + 2^x* x ln^2(2) + 2^(x +1) ln 2 --------------------------------------------------------- 9 cos (3x)

OpenStudy (welshfella):

- that was a long winded one I'll check it out on wolfram

OpenStudy (anonymous):

Thanks .the answer given in the book is (2/9) log(2/3)

OpenStudy (welshfella):

yes that looks good now plug in x = 0 for your limit

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

yes plugging in 0 the 2 middle terms = 0 and denominator = 9 so we have -2 ln3 + 2 ln2 ------------ 9 == 2ln 2 - 2 ln3 --------- 9 = (2/9) ln (2/3)

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

that was a tricky differentiation

OpenStudy (anonymous):

God bless !!! Thanks a ton both you'll guys !!!!!

OpenStudy (welshfella):

yw